OpenStudy (anonymous):

I am having a hard time solving for y in the following equation y-10=1/2(x-6)

OpenStudy (amistre64):

start by multipling the whole thing by 2 perhaps?

OpenStudy (amistre64):

is that (1/2) * (x-6) ?? or is it all part of a denominator?

OpenStudy (anonymous):

im not sure?

OpenStudy (anonymous):

Okay. SO what you have here is this: y - 10 = (x-6)/2 First off, add 10 to both sides. Then you have: y = (x-6)/2 + 10 Multiply both sides by 2 to eliminate the denominator. 2y = x-6+20 Which gives: 2y = x + 26 y = x/2 + 13 That is your final answer. If this helps, please give me a medal. http:/www.tutorsean.net

OpenStudy (anonymous):

is x/2 the same as 1/2x?